A continuous random variable x has the pdf given by

f(x)={ e-x for x>0

{0,ellsewhere

Find the mean and the standard deviation of

a)x

b)y=e(3/4)x

1
Expert's answer
2021-07-19T07:22:21-0400

a)


"E[X]=\displaystyle\int_{0}^{\infin}xe^{-x}dx"

"\int xe^{-x}dx"

"\int udv=uv-\int vdu"

"u=x, du=dx"

"dv=e^{-x}dx, v=\int e^{-x}dx=-e^{-x}"

"\int xe^{-x}dx=-xe^{-x}+\int e^{-x}dx=-xe^{-x}-e^{-x}+C_1"


"E[X]=\displaystyle\int_{0}^{\infin}xe^{-x}dx=\lim\limits_{t\to\infin}\displaystyle\int_{0}^{t}xe^{-x}dx"

"=\lim\limits_{t\to\infin}[-xe^{-x}-e^{-x}]\begin{matrix}
 t \\
 0
\end{matrix}=-0-0+0+1=1"


"E[X^2]=\displaystyle\int_{0}^{\infin}x^2e^{-x}dx"

"\int x^2e^{-x}dx"

"\int udv=uv-\int vdu"

"u=x^2, du=2xdx"

"dv=e^{-x}dx, v=\int e^{-x}dx=-e^{-x}"

"\int x^2e^{-x}dx=-x^2e^{-x}+2\int xe^{-x}dx"

"=-x^2e^{-x}-2xe^{-x}-2e^{-x}+C_2"



"E[X^2]=\displaystyle\int_{0}^{\infin}x^2e^{-x}dx=\lim\limits_{t\to\infin}\displaystyle\int_{0}^{t}x^2e^{-x}dx"

"=\lim\limits_{t\to\infin}[-x^2e^{-x}-2xe^{-x}-2e^{-x}]\begin{matrix}
 t \\
 0
\end{matrix}"

"=-0-0-0+0+0+2=2"

"Var(X)=\sigma^2=E[X^2]-(E[X])^2=2-(1)^2=1"

"\sigma=\sqrt{\sigma^2}=\sqrt{2}"


b)


"E[Y]=\displaystyle\int_{0}^{\infin}e^{(3/4)x}e^{-x}dx=\lim\limits_{t\to\infin}\displaystyle\int_{0}^{t}e^{(-1/4)x}dx""=\lim\limits_{t\to\infin}[-4e^{(-1/4)x}]\begin{matrix}
 t \\
 0
\end{matrix}=-0+4=4"

"E[Y^2]=\displaystyle\int_{0}^{\infin}e^{2(3/4)x}e^{-x}dx=\lim\limits_{t\to\infin}\displaystyle\int_{0}^{t}e^{(1/2)x}dx"

"=\lim\limits_{t\to\infin}[2e^{(1/2)x}]\begin{matrix}
 t \\
 0
\end{matrix}=\infin"


"Var(Y)=\sigma^2=E[Y^2]-(E[Y])^2=\infin"


"\sigma=\sqrt{\sigma^2}=\infin"